# On Strongly Regular Graphs and the Friendship Theorem

### v1

@@ 4 Spanning Subgraphs of Strongly Regular Graphs
A spanning subgraph is a subgraph obtained by removing some edges from the original graph while preserving all its vertices. Suppose that $\mathsf{G}$ and $\mathsf{H}$ are graphs of the same order, belonging to two given families of strongly regular graphs $\operatorname{\mathsf{srg}}(n,d,\lambda,\mu)$ and $\operatorname{\mathsf{srg}}(n,d^{\prime},\lambda^{\prime},\mu^{\prime})$, respectively (i.e., both graphs have identical orders). The specific structure of these graphs may not be necessarily specified (and recall that there may be various nonisomorphic strongly regular graphs within such a family [3, 2]). Next, we derive a necessary condition for one of these graphs to be a spanning subgraph of the other, relying on the Lovász $\vartheta$-functions of these graphs and their complements.
We next rely on the following simple lemma that relates between the Lovász $\vartheta$-functions of two graphs (not necessarily strongly regular), where one is a spanning subgraph of the other.
@@
Let $\mathsf{G}$ and $\mathsf{H}$ be strongly regular graphs on $n$ vertices, within the families $\operatorname{\mathsf{srg}}(n,d,\lambda,\mu)$ and $\operatorname{\mathsf{srg}}(n,d^{\prime},\lambda^{\prime},\mu^{\prime})$, respectively. Then, a necessary condition for $\mathsf{H}$ to be a spanning subgraph of $\mathsf{G}$ is that their parameters satisfy the inequality
@@
The result is obtained by combining Corollary 2.1 (see (9), (10)) and Lemma 4.1 (see (14)), followed by straightforward algebra. ∎
@@
Let $\mathsf{G}$ and $\mathsf{H}$ be two strongly regular graphs within the families $\operatorname{\mathsf{srg}}(45,28,15,21)$ and $\operatorname{\mathsf{srg}}(45,22,10,11)$, respectively (by [3], there exist strongly regular graphs with these parameters). The three inequalities in (16) are satisfied, so the question if $\mathsf{H}$ can be a spanning subgraph of $\mathsf{G}$ is left open according to these inequalities. However, by Corollary 4.1, $\mathsf{H}$ cannot be a spanning subgraph of $\mathsf{G}$ since the left-hand side of (17) is equal to $\frac{3\sqrt{5}+1}{11}=0.7007\ldots<1$, so the necessary condition in (17) is violated.
@@ 5 Induced Subgraphs of Strongly Regular Graphs
In continuation to Section 4, which is focused on spanning subgraphs of strongly regular graphs, this section derives necessary conditions for a graph to be an induced subgraph of a strongly regular graph belonging to a given family $\operatorname{\mathsf{srg}}(n,d,\lambda,\mu)$ with specified parameters (recall that an induced subgraph is obtained by deleting vertices from the original graph, together with all their incident edges). The analysis here relies not only on the Lovász $\vartheta$-function of the strongly regular graph but also on its graph energy.
The graph energy is a spectral graph invariant originally introduced in [7], where its application in chemistry was explored. A comprehensive study of graph energy can be found in the book [10].
@@
The energy of a graph $\mathsf{G}$, denoted by $\operatorname{\mathcal{E}}(\mathsf{G})$, is defined to be the sum of the absolute values of the eigenvalues of the adjacency matrix of $\mathsf{G}$, i.e.,
@@
The following lemma relates the energy and the Lovász $\vartheta$-function of an induced subgraph to those of the original graph. It is general and does not assume strong regularity, and we later apply it to strongly regular graphs. It relies on the Cauchy interlacing theorem, which is cited as follows:
@@
Inequality (20) follows from the Cauchy interlacing theorem applied to the adjacency matrix of the graph $\mathsf{G}$. Note that the adjacency matrix of the induced subgraph $\mathsf{H}$ is obtained from the adjacency matrix of $\mathsf{G}$ by deleting the rows and columns that correspond to the deleted vertices in the subgraph $\mathsf{H}$. ∎
@@
[…]
We next rely on the spectral properties of strongly regular graphs, which are listed as follows, to derive a closed-form expression of their energy. Note that if $\mathsf{G}$ is a strongly regular graph that belongs to the family $\operatorname{\mathsf{srg}}(n,d,\lambda,\mu)$, then $\mathsf{G}$ is connected if and only if $\mu>0$. In that case, the diameter of $\mathsf{G}$ is equal to 2 (a complete graph is excluded from the family of strongly regular graphs).
Theorem 5.2 (The adjacency spactra of strongly regular graphs, [2]).
The following properties hold:
@@
    A strongly regular graph has three distinct eigenvalues.
@@
    Let $\mathsf{G}$ be a connected strongly regular graph that belongs to the family $\operatorname{\mathsf{srg}}(n,d,\lambda,\mu)$. Then, the largest eigenvalue of its adjacency matrix is $\lambda_{1}(\mathsf{G})=d$ with multiplicity 1, and the other two distinct eigenvalues of its adjacency matrix are given by
@@
    with the respective multiplicities in (13).
@@
    Strongly regular graphs for which $2d+(n-1)(\lambda-\mu)\neq 0$ have integral eigenvalues and their respective multiplicities are distinct.
@@
By Theorem 5.2, the spectrum of a strongly regular graph has exactly three distinct eigenvalues. A closed-form expression for the energy of strongly regular graphs with general parameters does not seem, however, to appear in the literature (see, e.g., [2, 8, 9, 10]), and it is next introduced for the continuation of the analysis in this section.
@@
Let $\mathsf{G}$ be a strongly regular graph that belongs to the family $\operatorname{\mathsf{srg}}(n,d,\lambda,\mu)$ with $\mu>0$. Then, its energy is given by
@@
By the spectral properties of strongly regular graphs, the adjacency matrix of a strongly regular graph has three distinct eigenvalues. The largest eigenvalue of $\mathsf{G}$ is equal to $d$, the second largest eigenvalue of $\mathsf{G}$ is $\lambda_{2}(\mathsf{G})=p_{1}$ with multiplicity $m_{1}$, and the smallest eigenvalue of $\mathsf{G}$ is $\lambda_{\min}(\mathsf{G})=p_{2}$ with multiplicity $m_{2}$ (see (21) and (13)). Consequently, the energy of $\mathsf{G}$ is given by
$\operatorname{\mathcal{E}}(\mathsf{G})$ $=d+m_{1}|p_{1}|+m_{2}|p_{2}|$
$=d+\tfrac{1}{2}\,\Biggl{(}\lambda-\mu+\sqrt{(\lambda-\mu)^{2}+4(d-\mu)}\,\Biggr{)}\cdot\tfrac{1}{2}\,\Biggl{(}n-1-\frac{2d+(n-1)(\lambda-\mu)}{\sqrt{(\lambda-\mu)^{2}+4(d-\mu)}}\,\Biggr{)}$
@@
where the last equality in (23) can be verified by straightforward algebra, and relying on the connection between the four parameters of a strongly regular graph as given in Item 1 of Proposition 3.1. ∎
Example 5.1.
The adjacency spectrum of the Petersen graph is the multiset $\{3^{1},1^{5},(-2)^{4}\}$, where superscripts indicate the multiplicities of the eigenvalues. Hence, its graph energy is 16. Since it is a strongly regular graph with parameters $\operatorname{\mathsf{srg}}(10,3,0,1)$, their substitution into (22) confirms this result.
@@
with equality if $\mathsf{G}$ is a strongly regular graph with parameters
@@
Furthermore, strongly regular graphs with parameters as in (25) exist if
@@
Let $\mathsf{H}$ be an induced subgraph of a strongly regular graph $\mathsf{G}$ that belongs of the family $\operatorname{\mathsf{srg}}(n,d,\lambda,\mu)$ with $\mu>0$. Then, the following inequalities hold:
@@
    Specifically, if $\mathsf{H}$ belongs to the family $\operatorname{\mathsf{srg}}(n^{\prime},d^{\prime},\lambda^{\prime},\mu^{\prime})$, where $n^{\prime}\leq n$, $d^{\prime}\leq d$, $\lambda^{\prime}\leq\lambda$, and $\nu^{\prime}\leq\nu$, then
@@
Inequalities (26)–(29) follow from Lemma 5.1, combined with its specialization according to Corollary 2.1 and Lemma 5.2. ∎
@@
The family of triangular graphs consists of regular graphs obtained as the line graphs of complete graphs on at least three vertices. Let $T_{\ell}$ denote the line graph of the complete graph $\operatorname{\mathsf{C}}_{\ell}$, for all $\ell\geq 3$. This results in a strongly regular graph belonging to the family $\operatorname{\mathsf{srg}}(\tfrac{1}{2}\ell(\ell-1),2\ell-4,\ell-2,4)$ for all $\ell\geq 4$ (see [2, Section 1.1.7]). For $\ell\geq 4$ with $\ell\neq 8$, these connected, strongly regular graphs are unique for their respective parameters. Their uniqueness further implies that they are determined by their spectra (following from Theorem 34 of [16]).
Now, let $\mathsf{H}$ be a strongly regular graph that belongs to the family $\operatorname{\mathsf{srg}}(n^{\prime},d^{\prime},\lambda^{\prime},\mu^{\prime})$ for some specified (feasible) parameters. We consider whether $\mathsf{H}$ can appear as an induced subgraph of the triangular graph $T_{\ell}$, as a function of $\ell\geq 4$. Applying Item 2 of Corollary 5.1, if either inequality (28) or (29) is violated, then no strongly regular graph in the family $\operatorname{\mathsf{srg}}(n^{\prime},d^{\prime},\lambda^{\prime},\mu^{\prime})$ (including $\mathsf{H}$) can be an induced subgraph of $T_{\ell}$. This leads to a general conclusion for those values of $\ell$ where either inequality fails. Applying Item 2 of Corollary 5.1 to $T_{\ell}$ with the parameters:
$n=\tfrac{1}{2}\ell(\ell-1),\quad d=2(\ell-2),\quad\lambda=\ell-2,\quad\mu=4,$ (30)
yields, by (10),
$t=\ell-2,\quad\forall\,\ell\geq 4.$ (31)
With straightforward algebra, the inequalities (28) and (29) simplify to
[…]
@@
Consider the Gewirtz graph, denoted by $\mathsf{H}$, which is a connected, triangle-free, strongly regular graph belonging to the family $\operatorname{\mathsf{srg}}(56,10,0,2)$ (see [2, Section 10.20]). To clarify, since $\mu=2>0$, the graph $\mathsf{H}$ is connected, and $\lambda=0$ confirms that it is triangle-free. Uniqueness in this parameter family ensures it is determined by its spectrum [16]. Specializing (32) for $\mathsf{H}$, the first inequality is violated if and only if $\ell\leq 31$ (it dominates the second, energy-based, inequality in (32)). Hence, $\mathsf{H}$ cannot be an induced subgraph of $T_{\ell}$ for all $\ell\leq 31$. For $\ell\leq 11$, this result is trivial since $\tfrac{1}{2}\ell(\ell-1)<56$, meaning that the order of $T_{\ell}$ is too small to contain $\mathsf{H}$ as a subgraph. However, for $\ell\geq 32$, this question remains open, as both inequalities in (32) hold.

### v3

@@ 4 Spanning Subgraphs of Strongly Regular Graphs
A spanning subgraph is obtained by removing some edges from the original graph while preserving all its vertices. Let $\mathsf{G}$ and $\mathsf{H}$ be graphs of the same order, belonging to the strongly regular graph families $\operatorname{\mathsf{srg}}(n,d,\lambda,\mu)$ and $\operatorname{\mathsf{srg}}(n,d^{\prime},\lambda^{\prime},\mu^{\prime})$, respectively. The specific structure of these graphs may not necessarily be specified, as multiple nonisomorphic strongly regular graphs can exist within a given family [7, 9]. We next derive a necessary condition for one of these graphs to be a spanning subgraph of the other, relying on the Lovász $\vartheta$-functions of these graphs and their complements.
We rely on the following simple lemma, which relates the Lovász $\vartheta$-functions of two graphs (not necessarily strongly regular), where one is a spanning subgraph of the other.
@@
Let $\mathsf{G}$ and $\mathsf{H}$ be strongly regular graphs on $n$ vertices, belonging to the families $\operatorname{\mathsf{srg}}(n,d,\lambda,\mu)$ and $\operatorname{\mathsf{srg}}(n,d^{\prime},\lambda^{\prime},\mu^{\prime})$, respectively. Then, a necessary condition for $\mathsf{H}$ to be a spanning subgraph of $\mathsf{G}$ is that their parameters satisfy the inequality
@@
The result is obtained by combining Theorem 2.2 (see (10), (11)) and Lemma 4.1 (see (38)), followed by straightforward algebra. ∎
@@
Let $\mathsf{G}$ and $\mathsf{H}$ be strongly regular graphs belonging to the families $\operatorname{\mathsf{srg}}(45,28,15,21)$ and $\operatorname{\mathsf{srg}}(45,22,10,11)$, respectively (by [7], there exist strongly regular graphs with these parameters). The inequalities in (40) are satisfied, so the question if $\mathsf{H}$ can be a spanning subgraph of $\mathsf{G}$ is left open according to these inequalities. However, by Corollary 4.1, $\mathsf{H}$ (or any graph isomorphic to $\mathsf{H}$) cannot be a spanning subgraph of $\mathsf{G}$ since the left-hand side of (41) is equal to $\frac{3\sqrt{5}+1}{11}=0.7007\ldots<1$, so the necessary condition in (41) is violated.
The following result extends Corollary 4.1 to a more general setting, allowing for regular spanning subgraphs of a strongly regular graph and, even more broadly, regular spanning subgraphs of regular graphs. This extension relaxes the strong regularity requirement in Corollary 4.1, replacing it with the milder condition of regularity. This generalization comes, however, at a certain cost: it necessitates knowledge of the second-largest and least eigenvalues of their adjacency spectra, whereas Corollary 4.1 requires to only know the parameter vectors that characterize the strongly regular graphs $\mathsf{G}$ and $\mathsf{H}$. This distinction arises because nonisomorphic strongly regular graphs in the same family $\operatorname{\mathsf{srg}}(n,d,\lambda,\mu)$, for any feasible parameters $(n,d,\lambda,\mu)$, are cospectral, with their adjacency spectra uniquely determined by these parameters. In contrast, nonisomorphic $d$-regular graphs on $n$ vertices are not necessarily cospectral.
Corollary 4.2.
Let $\mathsf{G}$ and $\mathsf{H}$ be noncomplete and nonempty $d$-regular and $d^{\prime}$-regular graphs on $n$ vertices, with $d^{\prime}\leq d$. Then, $\mathsf{H}$ is a spanning subgraph of $\mathsf{G}$ if the following inequality is satisfied:
$\frac{n\bigl{(}1+\lambda_{2}(\mathsf{G})\bigr{)}}{n-d+\lambda_{2}(\mathsf{G})}+\frac{d^{\prime}}{\lambda_{n}(\mathsf{H})}\geq 1,$ (42)
where $\lambda_{2}(\mathsf{G})$ is the second-largest eigenvalue of the adjacency matrix of $\mathsf{G}$, and $\lambda_{n}(\mathsf{H})$ is the smallest eigenvalue of the adjacency matrix of $\mathsf{H}$.
Proof.
The necessary condition in (42) holds by combining inequality (39) with the lower bound on $\vartheta(\mathsf{G})$ and the upper bound on $\vartheta(\mathsf{H})$ in the leftmost and rightmost inequalities in (7), respectively. ∎
Remark 4.2.
Combining inequality (38) with the lower bound on $\vartheta(\overline{\mathsf{H}})$ and the upper bound on $\vartheta(\overline{\mathsf{G}})$ in the leftmost and rightmost inequalities in (8), respectively, gives
$\frac{n-d+\lambda_{2}(\mathsf{G})}{1+\lambda_{2}(\mathsf{G})}\leq-\frac{n\lambda_{n}(\mathsf{H})}{d^{\prime}-\lambda_{n}(\mathsf{H})},$ (43)
which is equivalent to (42) (note that for a noncomplete regular graph $\mathsf{G}$, its second-largest eigenvalue satisfies $\lambda_{2}(\mathsf{G})>-1$, and $\lambda_{n}(\mathsf{H})<0$ for a nonempty graph $\mathsf{H}$). Corollary 4.2 consequently contains a single inequality since the two inequalities in Lemma 4.1, combined with (7) and (8), yield an identical inequality.
@@ 5 Induced Subgraphs of Strongly Regular Graphs
An induced subgraph is obtained by deleting vertices from the original graph, along with all their incident edges. In analogy to Section 4, which is focused on spanning subgraphs of strongly regular graphs, and an extension of that analysis to regular graphs and spanning subgraphs, this section examines induced subgraphs in a similar manner. Specifically, we first derive necessary conditions for a graph to be an induced subgraph of a strongly regular graph in the family $\operatorname{\mathsf{srg}}(n,d,\lambda,\mu)$. This analysis relies not only on the Lovász $\vartheta$-function for strongly regular graphs but also on their energy.
The graph energy is a graph invariant originally introduced in [15], exploring its application in chemistry. A comprehensive treatment of graph energy can be found in [16].
@@
The energy of a graph $\mathsf{G}$ on $n$ vertices, denoted by $\operatorname{\mathcal{E}}(\mathsf{G})$, is given by
@@
The following lemma relates the energy and the Lovász $\vartheta$-function of an induced subgraph to those of the original graph. It is general, and it is then applied to strongly regular graphs, and extended to regular graphs. It partially relies on the Cauchy interlacing theorem, which is cited as follows:
@@
Inequality (46) follows from the Cauchy interlacing theorem (Theorem 5.1), applied to the adjacency matrix of the graph $\mathsf{G}$. Note that the adjacency matrix of the induced subgraph $\mathsf{H}$ is obtained from the adjacency matrix of $\mathsf{G}$ by deleting the rows and columns that correspond to the deleted vertices from $\mathsf{G}$. ∎
@@
Inequality (46) states that the energy of an induced subgraph is at most the energy of the original graph. However, this result does not generally extend to spanning subgraphs (see p. 64 of [16]). A simple counterexample demonstrating that the energy of a spanning subgraph can exceed that of the original graph is given by considering $\mathsf{G}=\operatorname{\mathsf{C}}_{4}$ and removing an edge to obtain the path graph $\mathsf{H}=\operatorname{\mathsf{P}}_{4}$ (on 4 vertices) as a spanning subgraph of $\mathsf{G}$. The adjacency spectra of $\mathsf{G}$ and $\mathsf{H}$ are, respectively, given by $\{-2,0,0,2\}$ and $\bigl{\{}\tfrac{1}{2}(\sqrt{5}+1),\,\tfrac{1}{2}(\sqrt{5}-1),\,-\tfrac{1}{2}(\sqrt{5}-1),-\tfrac{1}{2}(\sqrt{5}+1)\bigr{\}}$, so it follows that $\operatorname{\mathcal{E}}(\mathsf{G})=4<2\sqrt{5}=\operatorname{\mathcal{E}}(\mathsf{H})$. On the other hand, a sufficient condition where the energy of a spanning subgraph does not exceed that of the original graph is given in Theorem 4.20 of [16]. This highlights that the energy of a spanning subgraph can be either greater or smaller than that of the original graph, which is why such a condition is not included in Section 4.
We next utilize the spectral properties of strongly regular graphs, listed in Theorem 5.2, to derive a closed-form expression for their energy.
Theorem 5.2 (The adjacency spectra of strongly regular graphs, Chapter 21 of [11]).
The following spectral properties of strongly regular graphs hold:
@@
    A strongly regular graph has at most three distinct eigenvalues.
@@
    Let $\mathsf{G}$ be a connected strongly regular graph in the family $\operatorname{\mathsf{srg}}(n,d,\lambda,\mu)$ (i.e., $\mu>0$). Then, its adjacency spectrum consists of three distinct eigenvalues, where the largest eigenvalue is given by $\lambda_{1}(\mathsf{G})=d$ with multiplicity 1, and the other two distinct eigenvalues of its adjacency matrix are given by
@@
    with the respective multiplicities
@@
    Connected strongly regular graphs, for which $2d+(n-1)(\lambda-\mu)\neq 0$, have integral eigenvalues and their respective multiplicities are distinct.
@@
  6. (6)
    A disconnected strongly regular graph is a disjoint union of $m$ identical complete graphs $\operatorname{\mathsf{K}}_{r}$, where $m\geq 2$ and $r\in\mathbb{N}$. It belongs to the family $\operatorname{\mathsf{srg}}(mr,r-1,r-2,0)$, and its adjacency spectrum is $\{(r-1)^{[m]},(-1)^{[m(r-1)]}\}$, where superscripts indicate the multiplicities of the eigenvalues, thus having two distinct eigenvalues.
A closed-form expression for the energy of strongly regular graphs as a function of a general parameter vector $(n,d,\lambda,\mu)$ does not appear to be explicitly available in the literature (see, e.g., [9, 18, 19, 16]). Such an expression is next introduced for further analysis.
@@
The energy of a strongly regular graph $\mathsf{G}$ in the family $\operatorname{\mathsf{srg}}(n,d,\lambda,\mu)$ is given by
@@
A connected strongly regular graph $\mathsf{G}$ has an adjacency spectrum consisting of exactly three distinct eigenvalues, as determined in Item 2 of Theorem 5.2. The largest eigenvalue of $\mathsf{G}$ is $d$ with multiplicity 1, the second largest eigenvalue is $\lambda_{2}(\mathsf{G})=p_{1}$ with multiplicity $m_{1}$, and the smallest eigenvalue is $\lambda_{\min}(\mathsf{G})=p_{2}$ with multiplicity $m_{2}$ (see (47) and (48)). Consequently, by (44), the energy of $\mathsf{G}$ is given by
$\operatorname{\mathcal{E}}(\mathsf{G})=d+m_{1}|p_{1}|+m_{2}|p_{2}|.$ (50)
Substituting (47) and (48) into (50) gives
$\operatorname{\mathcal{E}}(\mathsf{G})$ $=d+\tfrac{1}{2}\,\Biggl{(}\lambda-\mu+\sqrt{(\lambda-\mu)^{2}+4(d-\mu)}\,\Biggr{)}\cdot\tfrac{1}{2}\,\Biggl{(}n-1-\frac{2d+(n-1)(\lambda-\mu)}{\sqrt{(\lambda-\mu)^{2}+4(d-\mu)}}\,\Biggr{)}$
@@
$=d+\tfrac{1}{2}(n-1)\sqrt{(\lambda-\mu)^{2}+4(d-\mu)}+\frac{2d+(n-1)(\lambda-\mu)}{\sqrt{(\lambda-\mu)^{2}+4(d-\mu)}}\cdot\tfrac{1}{2}(\mu-\lambda)$
$=d+\frac{2(n-1)(d-\mu)+d(\mu-\lambda)}{\sqrt{(\lambda-\mu)^{2}+4(d-\mu)}}$
$=d+\frac{\bigl{[}2(n-1)+\mu-\lambda\bigr{]}\,d-2(n-1)\mu}{\sqrt{(\lambda-\mu)^{2}+4(d-\mu)}}$
@@
where the last equality in (51) can be verified by relying on the connection between the four parameters of a strongly regular graph as given in Item 1 of Proposition 3.1, followed by straightforward algebra. This establishes (49) for all connected strongly regular graphs.
If $\mathsf{G}$ is a disconnected strongly regular graph then, by Item 6 of Theorem 5.2, it is a disjoint union of $m$ identical complete graphs $\operatorname{\mathsf{K}}_{r}$, where $m\geq 2$ and $r\in\mathbb{N}$. The adjacency spectrum of $\mathsf{G}$ is then given by $\{(r-1)^{[m]},(-1)^{[m(r-1)]}\}$, so the graph energy is $\operatorname{\mathcal{E}}(\mathsf{G})=2m(r-1)$. That strongly regular graph $\mathsf{G}$ belongs to the family $\operatorname{\mathsf{srg}}(mr,r-1,r-2,0)$, and substituting the parameters $n=mr$, $d=r-1$, $\lambda=r-2$, and $\mu=0$ into (49) verifies that the latter equality extends to disconnected strongly regular graphs. ∎
@@
with equality if $\mathsf{G}$ is a strongly regular graph in the family
@@
Furthermore, such strongly regular graphs exist if one of the following conditions hold:
@@
Let $\mathsf{H}$ be an induced subgraph of a strongly regular graph $\mathsf{G}$ belonging to the family $\operatorname{\mathsf{srg}}(n,d,\lambda,\mu)$. Then, the following inequalities hold:
@@
    Specifically, if $\mathsf{H}$ is in the family $\operatorname{\mathsf{srg}}(n^{\prime},d^{\prime},\lambda^{\prime},\mu^{\prime})$, where $n^{\prime}\leq n$, $d^{\prime}\leq d$, $\lambda^{\prime}\leq\lambda$, and $\nu^{\prime}\leq\nu$, then
@@
Inequalities (54)–(57) follow from Lemma 5.1, combined with its specialization to strongly regular graphs according to Theorem 2.2 and Lemma 5.2. ∎
Remark 5.2.
The computational complexity of the Lovász $\vartheta$-function for a graph $\mathsf{G}$ on $n$ vertices, obtained by numerically solving the SDP problem in (3), is polynomial in $n$ and $\log\frac{1}{\varepsilon}$, where $\varepsilon$ denotes the computation precision (see Theorem 11.11 of [6]). The constrained convex optimization problem in (3) involves $\tfrac{1}{2}n(n+1)$ optimization variables – specifically, the entries of the $n\times n$ positive semidefinite matrix ${\bf{B}}$ – and at most quadratically many equality constraints in $n$ (scaling linearly in $n$ for sparse graphs). Consequently, while the Lovász $\vartheta$-function remains computationally feasible for graphs with up to several hundred vertices, it becomes impractical for significantly larger instances. Checking the satisfiability of the necessary conditions for spanning or induced subgraphs in Lemmata 4.1 and 5.1 is thus infeasible for large $n$. However, computing only the second-largest or smallest eigenvalue of the adjacency matrix can be efficiently performed using iterative methods, which avoid computing the entire spectrum. These methods are particularly well-suited for large graphs, enabling the verification of the necessary conditions for spanning or induced subgraphs of regular graphs in Corollary 4.2 and Item 1 of Corollary 5.1, even for large instances. Furthermore, verifying the necessary conditions in Corollary 4.1 and Item 2 of Corollary 5.1 for spanning or induced strongly regular subgraphs of strongly regular graphs is computationally straightforward and imposes no practical limitations, even for exceedingly large strongly regular graphs.
Example 5.1 (Cycles as induced subgraphs of the Shrikhande Graph).
The Shrikhande graph, denoted by $\mathsf{G}$, is a strongly regular graph in the family $\operatorname{\mathsf{srg}}(16,6,2,2)$ (there are two such nonisomorphic strongly regular graphs, see Section 10.6 of [9]). The graph $\mathsf{G}$ is thus a strongly regular graph in which every pair of vertices has exactly two common neighbors (by the friendship theorem, it stays in contrast to the non-existence of a regular graph, except for a triangle, in which every pair of vertices has exactly one common neighbor). It can be verified that $\mathsf{G}$ contains induced cycles of length 3, 4, 5, 6, and 8, but no induced cycles of length 7 or greater than 8. Figure 2 illustrates two such induced cycles, with a cycle of length 6 shown on the left plot and a cycle of length 8 on the right plot, both generated using the SageMath software [13].
Figure 2: The Shrikhande graph $\mathsf{G}$ and two induced cycles of length 6 (left plot) and 8 (right plot), with their edges highlighted in red.
Let $\mathsf{H}=\operatorname{\mathsf{C}}_{\ell}$ denote a cycle of length $\ell\geq 3$. The necessary conditions in (54) and (55) for a graph isomorphic to $\operatorname{\mathsf{C}}_{\ell}$ to be an induced subgraph of $\mathsf{G}$ are given by $\vartheta(\operatorname{\mathsf{C}}_{\ell})\leq 4$ and $\operatorname{\mathcal{E}}(\operatorname{\mathsf{C}}_{\ell})\leq 36$, respectively. These follow from the equalities $\vartheta(\mathsf{G})=4$ by Theorem 2.2 and $\operatorname{\mathcal{E}}(\mathsf{G})=36$ by Lemma 5.2. It can be verified that the former condition on the Lovász $\vartheta$-function is more restrictive in the studied case than the latter condition on the graph energy, so we focus on the necessary condition $\vartheta(\operatorname{\mathsf{C}}_{\ell})\leq 4$. By [5], for all integers $\ell\geq 3$,
[…]
Consequently, by (58), for all $\ell>8$, we get $\vartheta(\operatorname{\mathsf{C}}_{\ell})\geq\vartheta(\operatorname{\mathsf{C}}_{9})=4.361\ldots>4=\vartheta(\mathsf{G})$, verifying that $\mathsf{G}$ does not contain any induced cycle of length $\ell>8$. Furthermore, $\vartheta(\operatorname{\mathsf{C}}_{8})=4=\vartheta(\mathsf{G})$ and $\mathsf{G}$ contains a cycle of length 8 as an induced subgraph (see the right plot of Figure 2). This provides an example where the necessary condition holds with equality, coinciding with the existence of the longest possible induced cycle. However, since $\mathsf{G}$ does not contain an induced cycle of length 7, despite $\vartheta(\operatorname{\mathsf{C}}_{7})=3.31766\ldots<4=\vartheta(\mathsf{G})$, this serves as a counterexample where the necessary conditions in Item 1 of Corollary 5.1 fail to exclude the existence of an induced cycle of length 7 in $\mathsf{G}$.
Another counterexample where the necessary conditions in Item 1 of Corollary 5.1 fail to exclude the existence of an induced cycle of a given length is a connected, triangle-free graph $\mathsf{G}$ of energy at least 4 (noting that $\operatorname{\mathcal{E}}(\operatorname{\mathsf{C}}_{3})=4$). By definition, such a graph does not contain any triangle as a subgraph, yet the considered necessary conditions do not rule out the absence of an induced triangle in $\mathsf{G}$. Among the connected strongly regular graphs, there are seven currently known triangle-free graphs, and all of them are determined by their adjacency spectrum (see, e.g., Section 4.7 of [20]).
@@
Consider the Gewirtz graph, denoted by $\mathsf{H}$, which is a connected, triangle-free, strongly regular graph in the family $\operatorname{\mathsf{srg}}(56,10,0,2)$ (see Section 10.20 of [9]). To clarify, since $\mu=2>0$, the graph $\mathsf{H}$ is connected, and $\lambda=0$ confirms that it is triangle-free. Uniqueness in this parameter family ensures it is determined by its spectrum [20]. Specializing (LABEL:eq1:_19.02.25) for $\mathsf{H}$, the first inequality is violated if and only if $\ell\leq 31$ (it dominates the second, energy-based, inequality in (LABEL:eq1:_19.02.25)). Hence, $\mathsf{H}$ cannot be an induced subgraph of $T_{\ell}$ for all $\ell\leq 31$. For $\ell\leq 11$, this result is trivial since $\tfrac{1}{2}\ell(\ell-1)<56$, meaning that the order of $T_{\ell}$ is too small to contain $\mathsf{H}$ as a subgraph. However, for $\ell\geq 32$, this question remains open, as both inequalities in (LABEL:eq1:_19.02.25) hold.